OpenStudy (anonymous):

Please solve, 6x^4 - 11x^2 + 3 = 0

OpenStudy (anonymous):

http://www.myalgebra.com/algebra_solver.aspx plug into this site

myininaya (myininaya):

6*3=18 (-6)(-3)=18 (-2)(-9)=18 -2-9=11 6x^4-2x^2-9x^2+3=0 2x^2(3x^2-1)-3(3x^2-1)=0 (3x^2-1)(2x^2-3)=0 this should help

OpenStudy (anonymous):

(3x^2-1)(2x^2-3) easy

myininaya (myininaya):

set both =0 3x^2-1=0 and 2x^2-3=0